# QUSL: Quantum Unsupervised Image Similarity Learning with Enhanced Performance

### v2

@@ 4 Experiments and Result
This section presents a rigorous experimental evaluation of the QUSL model across diverse datasets, providing comprehensive empirical evidence to demonstrate its effectiveness, robustness, and superiority compared to state-of-the-art quantum methods.
@@ 4.1 Experiment Framework
In the domain of unsupervised image similarity tasks, the Flickr landscape, COCO, and DISC21 datasets have been selected for training and testing purposes. Flickr landscape [47], comprising unlabeled colorful images depicting diverse natural landscapes, serves to explore the performance of QUSL in image similarity detection and feature learning without explicit category labels. COCO [48], widely used for tasks like object detection and segmentation, offers a large-scale collection of labeled images with varied scenes and objects. By removing labels from COCO, we aim to scrutinize QUSL’s unsupervised learning capability further while assessing its adaptability to diverse and complex image datasets. DISC21 [49], explicitly designed for image similarity detection and utilized as a benchmark in the NeurIPS’21 Image Similarity Challenge (ISC2021), emphasizes image transformations, including automated conversion, manual editing, and ML-based operations. This dataset is instrumental in evaluating QUSL’s real-world applicability, particularly in social media contexts. The essential details of these datasets and the experimental configurations in our task are outlined in Table 1.
@@
COCO $80$ object categoriess $\leq 800\times 800\times 3^{b}$ $14$
DISC21 $21$ object categories $\leq 800\times 800\times 3$ $14$
@@
    landscape, COCO, and DISC21, do not utilize existing labels in the experiments.
  • b
    COCO and DISC21 feature a plethora of image dimensions; herein present the maximum dimensions.
The experimental setup was configured using Python3, incorporating a hybrid combination of the Qiskit and MindSpore [50] frameworks. Numerical simulations were executed on a workstation (CPU: Intel I9 9900k, GPU: GTX3090). Guided by pre-experiment instructions, an evolutionary algorithm tailored for quantum circuits was deployed, featuring a population size of $20$ and a maximum evolution generation of $20$. The gate set constituting the quantum circuit comprised ${R_{X},R_{Y},R_{Z},H,CNOT}$. Images within the training set underwent uniform processing to ensure consistent dimensions. The number of qubits utilized for each scenario is also delineated in Table  1. Following the selection of anchor images, random images from the training set were chosen to form $A,P,N$-triplets. During training, any potential labels in the dataset were disregarded to align with the unsupervised setting. For all models trained on similarity tasks, $1000$ similarity tests were conducted, and model performance was assessed using correlation detection.
@@ 4.2 Evaluation and Analysis
    Five independent training and testing sessions were conducted at each perturbation level, with correlation values reported to three decimal places.
@@
    $\pm$” represents the standard deviation, reflecting the dispersion of sample data, as it does elsewhere in this paper. A smaller value indicates higher stability of model performance.
@@
Controlled experiment were designed to comprehensively examine the performance efficacy and robustness of QUSL on three datasets (landscape, COCO, and DISC21), while concurrently benchmarking against SliQ. QUSL and SliQ underwent five training sessions and correlation tests of equal scale on each of the three datasets, recording the best results obtained during the training process, which are presented in Fig. 4 and Table. 3.
Figure 4: Performance evaluation of QUSL and SliQ across different datasets. In the violin plot, the blue and purple areas represent the distribution states of correlation obtained by SliQ and QUSL, respectively.
@@
Dataset landscape COCO DISC21
Correlationa SliQ QUSL(ours) SliQ QUSL(ours) SliQ QUSL(ours)
1st $0.835$ 0.889 $0.757$ 0.876 $0.758$ 0.861
2st $0.814$ $0.837$ $0.757$ $0.875$ $0.649$ $0.847$
3st $0.758$ $0.829$ $0.722$ $0.852$ $0.593$ $0.825$
4st $0.692$ $0.822$ $0.699$ $0.822$ $0.592$ $0.787$
5st $0.657$ $0.814$ $0.668$ $0.821$ $0.528$ $0.777$
Mean value $0.751$ (±$(\pm 0.078)$ 0.838 (±$(\pm\textbf{0.031})$ $0.721$ (±$(\pm 0.035)$ 0.849 (±$(\pm\textbf{0.027})$ $0.624$ (±$(\pm 0.096)$ 0.819 (±$(\pm\textbf{0.037})$
Optimization rate - $\uparrow 8.7\%$ $\uparrow 12.8\%$ $\uparrow 19.5\%$
@@
    The correlation obtained from multiple experiments is sorted in descending order.
In all three datasets, QUSL demonstrated significant enhancements in both performance and stability. In comparison to SliQ, QUSL exhibited a remarkable improvement in correlation, with increases ranging from a minimum of $8.7\%$ to a maximum of $19.5\%$. Regarding stability, QUSL displayed a distinct advantage over SliQ in terms of the standard deviation of performance means, particularly evident in the DISC21 dataset, where SliQ’s performance exhibited substantial fluctuations, while QUSL’s stability remained largely unaffected by the application scenarios.
QUSL’s superior performance and stability can be primarily attributed to its capacity to adapt to dataset characteristics through the evolutionary process. Conversely, the parameter adjustment of template-based variational quantum circuits may encounter difficulties in capturing the intrinsic properties of the dataset, resulting in performance degradation under equivalent training conditions. This set of experiments serves to highlight the transferability of QUSL across various domains.
@@
Critical quantum resource performance
Dataset landscape COCO DISC21
Correlation CNOT count Circuit depth CNOT count Circuit depth CNOT count Circuit depth
1 $35$ $36$ $30$ $33$ 17 $26$
2 $\textbf{20}^{a}$ 20 $33$ $40$ $24$ $30$
3 $26$ $28$ $36$ $44$ $25$ 25
4 $24$ $27$ 25 27 $36$ $42$
5 $25$ $26$ $25$ $31$ $26$ $32$
Mean value $26.0^{b}$(±$(\pm 4.9)$ $27.4$(±$(\pm 5.1)$ $29.8$(±$(\pm 4.2)$ $35.0$(±$(\pm 6.2)$ $25.6$(±$(\pm 6.3)$ $31.0$(±$(\pm 6.1)$
[…]
@@ 5 $25$ $26$ $25$ $31$ $26$ $32$
  • b
  • c
    The optimization rate refers to the decrease ratio of the average CNOT count and circuit depth compared to their respective values in SliQ, where the CNOT count is $52$ and the circuit depth is $64$.
Across the landscape, COCO, and DISC21 datasets, the quantum circuits derived through QUSL exhibit reductions of around $50\%$ in both CNOT counts and circuit depth. This indicates that within acceptable fluctuations, the quantum circuits generated by QUSL significantly outperform SliQ in key performance indicators, directly contributing to the model’s advantage in quantum resource utilization. The demonstration of high-performance quantum circuits is presented in A.
A set of validation experiments on quantum computers demonstrated the practical usability and significant advantages of QUSL in the NISQ era. Table 5 presents a comparison of the running times between the high-performance quantum circuits used in QUSL and the template circuits employed in SliQ.
@@
Optimization rate - $\downarrow\textbf{0.13}$ $\downarrow\textbf{0.11}$ $\downarrow\textbf{0.13}$
@@
Through five independent circuit runs on the three datasets, it can be observed that QUSL consistently reduces the running time of individual quantum circuits. The time advantage at the single-circuit level is significantly amplified by the large number of circuit runs during the training process, which is reflected in the model’s training efficiency and practical application efficiency. This advantage of QUSL becomes more evident in large-scale image similarity tasks.
Through numerical simulations and a series of experiments on quantum computers, the QUSL algorithm demonstrates robust high performance in image similarity tasks, achieving higher average correlations across multiple datasets and significantly improving stability on challenging datasets like DISC21. Detailed analysis of the training process reflects QUSL’s effective ability to capture visual similarities among images. Combined with validation on quantum computers, QUSL significantly reduces quantum circuit depth and CNOT counts compared to SliQ’s variational template. The aforementioned advantages directly translate into improved utilization of quantum resources in NISQ environments and achieve higher performance and efficiency in large-scale image similarity scenarios that are closer to real-world applications.
Overall, QUSL emerges as a high-performance method for image similarity detection, surpassing current state-of-the-art quantum image similarity detection methods in terms of stability, correlation, and quantum resource efficiency.
@@ 5 Discussion and Futher Work
Upon comparison with our work, SliQ’s [23] impressive performance showcases the advantage of quantum computing in image processing. In the current methodology, we have adopted a scheme inspired by the work of SliQ, which utilizes the correlation between the computed model loss and the Euclidean distance to evaluate the performance of the model. This ingenious and transferrable approach addresses the challenge of objectively assessing image similarity in unlabeled datasets, proving effective within the limited application scenarios addressed in our work. However, this evaluation method remains open to refinement. For instance, explore more suitable correspondences between losses and classical image similarity detection methods within different numerical ranges. This refinement process stands as a significant future endeavor within our research trajectory.
During the construction process of $A,P,N$ triplets, we employed noise to generate perturbed versions of reference images as positive instances within the triplets, and the efficacy of this strategy was validated through subsequent experimental analysis. While noise is conventionally perceived as an impediment in both classical and quantum contexts — noise cuts both ways — it can also serve as a valuable asset for specific tasks [51]. Although noise mitigation techniques have been extensively explored in quantum computing, acknowledging noise as a potentially beneficial factor paves the way for novel approaches and methodologies.
@@ 6 Conclusion
The QUSL method proposed in this work represents an advanced quantum approach tailored to address the challenges inherent in image similarity detection. A series of numerical simulations and quantum computer experiments reveal a substantial enhancement in similarity detection correlation compared to contemporary quantum methodologies, accompanied by significant reductions in quantum resource utilization. These empirical findings underscore the promising potential of the QUSL framework in effectively modeling similarity within expansive, unlabeled image datasets. Additionally, as a methodological framework, it exhibits transferability across diverse task scenarios. Further exploration of QUSL will revolve around more precise and objective similarity assessment and the utilization of noise for enhanced performance.
@@ Appendix A The quantum circuits obtained by QUSL
Figure 10: The optimal-performing quantum circuit obtained by QUSL on the landscape dataset exhibits a correlation performance of $0.889$, with $26$ CNOT operations and a circuit depth of $28$.
Figure 11: The optimal-performing quantum circuit obtained by QUSL on the DISC21 dataset exhibits a correlation performance of $0.861$, with $17$ CNOT operations and a circuit depth of $26$.
Figure 12: The optimal-performing quantum circuit obtained by QUSL on the COCO dataset exhibits a correlation performance of $0.876$, with $36$ CNOT operations and a circuit depth of $44$.
@@ Appendix B Quantum circuits with similar fitness but significantly different structures
Figure 13: Quantum circuits with significant structural differences but similar fitness levels during training. These circuits are derived from the training process on the COCO dataset, exhibiting fitness differences on the order of $10^{-4}$, yet their quantum circuit depths and CNOT counts differ by almost a factor of two.
Appendix C The model’s performance in classification tasks

### v4

@@ 4 Experiments and Result
This section presents an experimental evaluation of the QUSL model, demonstrating its effectiveness, robustness, and superiority compared to state-of-the-art quantum methods.
@@ 4.1 Experiment Framework
In the domain of unsupervised image similarity tasks, we have selected five datasets for training and testing purposes. Table 1 outlines the essential details of these datasets and the experimental configurations.
@@
COCO $80$ object categories $50\times 50\times 3$ $14$
DISC21 $21$ object categories $50\times 50\times 3$ $14$
CIFAR_10 $10$ object categories $32\times 32\times 3$ $14$
ImageNet $1000$ object categories $50\times 50\times 3$ $14$
@@
    All datasets do not utilize existing labels in the experiments.
The Flickr landscape Men et al. ((2022), comprising unlabeled colorful images depicting diverse natural landscapes, along with the label-removed COCO Lin et al. ((2014) and CIFAR__\__10 Krizhevsky et al. (2009), serve to initially examine QUSL’s adaptability to unsupervised tasks. Experiments on ImageNet Deng et al. (2009) further test QUSL’s performance. DISC21 Douze et al. (2021), specifically designed for image similarity detection tasks, is instrumental in evaluating QUSL’s applicability in real-world scenarios, particularly in social media contexts.
The experimental setup was configured using Python3, incorporating a hybrid combination of the Qiskit and MindSpore Developer (2021) frameworks. Numerical simulations were executed on a workstation (CPU: Intel I9 9900k, GPU: GTX3090). Guided by pre-experiment instructions, the evolutionary algorithm featured a population size of $20$ and a maximum evolution generation of $20$. The gate set constituting the quantum circuit comprised ${R_{X},R_{Y},R_{Z},H,CNOT}$. Images within the training set underwent uniform processing to ensure consistent dimensions. Following the selection of anchor images, random images from the training set were chosen to form $A,P,N$-triplets. For all models, $1000$ similarity tests were conducted, and model performance was assessed using correlation detection.
@@ 4.2 Evaluation and Analysis
    Five independent training and testing sessions were conducted at each level.
@@
    $\pm$” represents the standard deviation, reflecting the dispersion of sample data, as it does elsewhere in this paper.
@@
Controlled experiment were designed to comprehensively examine the performance efficacy and robustness of QUSL on five datasets, while concurrently benchmarking against SliQ. QUSL and SliQ underwent five training sessions and correlation tests of equal scale on each of the five datasets, recording the best results obtained during the training process, which are presented in Fig. 5 and Table. 3.
Figure 5: Performance evaluation of QUSL and SliQ across different datasets. In the violin plot, the blue and purple areas represent the distribution states of correlation obtained by SliQ and QUSL, respectively. Error bars represent $95\%$ confidence intervals
@@
Dataset Model 1st 2nd 3rd 4th 5th Mean value Improva.
landscape SliQ 0.835 0.814 0.758 0.692 0.657 0.751 ($\pm$0.078) $\uparrow$8.7%
QUSL 0.889 0.837 0.829 0.822 0.814 0.838 ($\pm$0.031)
COCO SliQ 0.757 0.757 0.722 0.699 0.668 0.721 ($\pm$0.035) $\uparrow$12.8%
QUSL 0.876 0.875 0.852 0.822 0.821 0.849 ($\pm$0.027)
DISC21 SliQ 0.758 0.649 0.593 0.592 0.528 0.624 ($\pm$0.096) $\uparrow$19.5%
QUSL 0.861 0.847 0.825 0.787 0.777 0.819 ($\pm$0.037)
CIFAR_10 SliQ 0.743 0.742 0.742 0.686 0.619 0.706 ($\pm$0.052) $\uparrow$11.9%
QUSL 0.890 0.792 0.769 0.766 0.731 0.790 ($\pm$0.059)
Imagnet SliQ 0.761 0.706 0.681 0.647 0.679 0.695 ($\pm$0.042) $\uparrow$17.0%
QUSL 0.848 0.824 0.818 0.809 0.765 0.813 ($\pm$0.030)
@@
    Improvement of QUSL(ours) over SliQ in terms of mean correlation performance.
In all five datasets, QUSL demonstrated significant enhancements in both performance and stability. In comparison to SliQ, QUSL exhibited a remarkable improvement in correlation, with increases ranging from a minimum of $8.7\%$ to a maximum of $19.5\%$. Regarding stability, QUSL displayed a distinct advantage over SliQ in terms of the standard deviation of performance means, particularly evident in the DISC21 dataset, where SliQ’s performance exhibited substantial fluctuations, while QUSL’s stability remained largely unaffected by the application scenarios.
QUSL’s performance improvements and high robustness can be primarily attributed to its ability to adapt to dataset characteristics through evolutionary algorithms. Conversely, the parameter adjustment of template-based variational quantum circuits may encounter difficulties in capturing the intrinsic properties of the dataset, resulting in performance degradation under equivalent training conditions. This set of experiments directly demonstrates QUSL’s performance in scenarios close to real-world applications and showcases its robustness across different types of image tasks.
@@
Dataset Metric 1st 2nd 3rd 4th 5th Mean valuea Improvb
landscape CNOT count 35 20c 26 24 25 26.0 ($\pm$4.9) $\downarrow$50.0%
Circuit depth 36 20 28 27 26 27.4 ($\pm$5.1) $\downarrow$57.2%
COCO CNOT count 30 33 36 25 25 29.8 ($\pm$4.2) $\downarrow$42.7%
Circuit depth 33 40 44 27 31 35.0 ($\pm$6.2) $\downarrow$45.3%
DISC21 CNOT count 17 24 25 36 26 25.6 ($\pm$6.3) $\downarrow$50.8%
Circuit depth 26 30 25 42 32 31.0 ($\pm$6.1) $\downarrow$51.6%
CIFAR_10 CNOT count 32 19 28 18 36 26.6 ($\pm$7.1) $\downarrow$48.8%
Circuit depth 36 23 34 27 29 29.8 ($\pm$4.7) $\downarrow$53.4%
Imagnet CNOT count 41 12 23 32 26 26.8 ($\pm$9.7) $\downarrow$48.5%
Circuit depth 45 16 31 34 35 32.2 ($\pm$9.3) $\downarrow$49.7%
@@
  • b
    Improvement of QUSL(ours) over SliQ in terms of the decrease ratio of the average CNOT count and circuit depth compared to their respective values in SliQ.
  • c
@@
Across all datasets used, the quantum circuits derived through QUSL exhibit reductions of around $50\%$ in both CNOT counts and circuit depth. This indicates that within acceptable fluctuations, the quantum circuits generated by QUSL significantly outperform SliQ in key performance indicators, directly contributing to the model’s advantage in quantum resource utilization. The demonstration of high-performance quantum circuits is presented in A.
Validation experiments on quantum computers were designed to examine the potential advantages of QUSL in the NISQ era. The experiments were conducted on the $ibm_{o}saka$ quantum computer with $500$ shots. Table 5 presents a comparison of the running times between the high-performance quantum circuits used in QUSL and the template circuits employed in SliQ.
@@
Optimization - $\downarrow\textbf{0.13}$ $\downarrow\textbf{0.11}$ $\downarrow\textbf{0.13}$
@@
Through five independent circuit runs, it can be observed that QUSL consistently reduces the running time of individual quantum circuits. The time advantage at the single-circuit level is significantly amplified by the large number of circuit runs during the training process, which will be reflected in the model’s training efficiency and practical application efficiency. This advantage of QUSL becomes more evident in large-scale tasks.
Through numerical simulations and a series of experiments on quantum computers, the QUSL algorithm demonstrates robust performance improvements in image similarity tasks, achieving higher average correlations across multiple datasets. In challenging datasets closer to real-world scenarios, such as DISC21, QUSL exhibits higher performance and stronger stability. More importantly, compared to SliQ’s variational template, QUSL significantly reduces quantum resource consumption. Combined with validation on quantum computers, QUSL demonstrates higher learning efficiency in NISQ environments relative to SliQ.
Overall, QUSL, as an unsupervised quantum image similarity model, surpasses current state-of-the-art quantum image similarity detection methods in terms of accuracy, robustness, and quantum resource efficiency.
@@ 5 Discussion and Futher Work
Along with our work, SliQ’s Silver et al. ((2023) performance demonstrates that quantum computing can be applied to image processing tasks with potential efficiency gains. Due to NISQ limitations, quantum machine learning algorithms native to quantum computers are challenging to implement. Both QUSL and SliQ represent quantum transfers of classical unsupervised image tasks, particularly contrastive learning tasks. In comparison, QUSL goes further than the latest methods in realizing a quantum implementation of the classical triplet method. Furthermore, this work reveals the importance of quantum circuit design in quantum unsupervised learning tasks. As an efficient method, heuristic quantum circuit architecture search can design high-performance circuits that adapt to task requirements. This directly alleviates the dependence of complex unsupervised tasks on deep, parameter-rich PQCs, making it more suitable for the NISQ era. This work provides a preliminary exploration of large-scale unsupervised contrastive learning in the quantum domain.
These methods utilize the correlation between the computed model loss and classical similarity norm to evaluate model performance. This ingenious and transferable approach addresses the challenge of objectively assessing image similarity in unlabeled datasets, proving effective within the limited application scenarios addressed in our work. However, this evaluation method remains open to refinement. For instance, exploring more suitable correspondences between losses and classical image similarity detection methods within different numerical ranges. This refinement process stands as a significant future endeavor within our research trajectory.
In the process of constructing A,P,N-triplets, noise is used for data augmentation. The efficacy of this strategy was validated through subsequent experimental analysis. While noise is conventionally perceived as an impediment in both classical and quantum contexts — noise cuts both ways — it can also serve as a valuable asset for specific tasks Liu et al. (2023). Although noise mitigation techniques have been extensively explored in quantum computing, acknowledging noise as a potentially beneficial factor paves the way for novel approaches and methodologies
@@ 6 Conclusion
The QUSL method proposed in this work focuses on quantum unsupervised image similarity detection tasks. QUSL employs heuristic methods to adapt quantum circuits to dataset characteristics and utilizes noise-based data augmentation for triplet construction. A series of numerical simulations and quantum computer experiments reveal that, compared to the current state-of-the-art quantum methods, QUSL further improves the accuracy and robustness of the task. QUSL achieves up to a $19.5\%$ increase in correlation and reduces the use of critical quantum resources by over $50\%$. As a model framework, QUSL demonstrates transferability applicable to other task scenarios. This work contributes to further integration of classical unsupervised learning techniques with quantum computing, showcasing the potential of quantum methods in complex image processing tasks.
@@ Appendix A The quantum circuits obtained by QUSL
The quantum circuits with the best expressive capabilities obtained by QUSL on landscape and DISC21 are shown below. Summarizing the optimal circuits obtained across all datasets reveals some interesting patterns and structures. $\{R_{X},R_{Y},R_{Z}\}$ and $CNOT$ gates present complex interweaving, forming an intricate high-dimensional Hilbert space. The $CNOT$-$R_{X}$-$R_{Y}$-$R_{Z}$ sequence appears frequently in multiple circuits, potentially serving as an effective module. Another prominent pattern is multiple qubits connected to a single qubit through $CNOT$ gates, which may effectively transform this qubit into an attention center through entanglement, thereby achieving crucial feature extraction.
Figure 11: The optimal-performing quantum circuit obtained by QUSL on the landscape, with $26$ CNOT operations and a circuit depth of $28$.
Figure 12: The optimal-performing quantum circuit obtained by QUSL on the DISC21, with $17$ CNOT operations and a circuit depth of $26$.
@@ Appendix B Quantum circuits with similar fitness but significantly different structures
During the training process, some circuits with similar expressive capabilities but enormous structural differences emerged, as shown in the figure below. These circuits are derived from the training process on the COCO dataset, exhibiting fitness differences on the order of $10^{-4}$, yet their quantum circuit depths and CNOT counts differ by almost a factor of two. This phenomenon poses challenges for population diversity.
Figure 13: Quantum circuits with significant structural differences but similar fitness levels.
Appendix C Performance Evaluation of QUSL in classification tasks
@@
Appendix D Performance Evaluation of QUSL in NISQ Environments
To validate QUSL’s potential for practical application in the NISQ era, we conducted simulated experiments using the quantum noise library provided by Pennylane. We added composite noise with a degree of $0.045$ (including bit flip, phase flip, and depolarizing channel noise) to the experiments. Additionally, to improve efficiency, we compressed the images through downsampling. Five independent trials were conducted in both clean and noisy environments with CIFAR_10:
Table 7: Performance of QUSL in noise-free and noisy environments
Environment 1st 2nd 3rd 4th 5th Average (±$(\pm std)$
Noise-free 0.8094 0.7941 0.7824 0.7778 0.7755 $0.7878(\pm 0.0143)$
Noisy 0.8063 0.7492 0.7246 0.7411 0.7067 $0.7226(\pm 0.0354)$
Results show that QUSL’s performance remained stable in noisy environments, with only a slight decrease in robustness. We speculate that the noise primarily affected the consistency in the triplet construction process, thereby influencing feature extraction. Nevertheless, QUSL’s excellent performance in noisy environments demonstrates its potential for application on actual NISQ devices.
Appendix E Performance Comparison of QUSL with Classical Methods
To evaluate QUSL’s performance, we designed a set of comparative experiments in CIFAR__\__10. The classical baseline used a medium-scale fully connected neural network with $4$ hidden layers and approximately $1$ million parameters, also employing the triplet strategy.
Table 8: Performance comparison of Classical Network, SliQ, and QUSL
Method 1st 2nd 3rd 4th 5th Average (±$(\pm std)$
Classical 0.833 0.771 0.766 0.745 0.742 $0.771(\pm 0.036)$
SliQ 0.743 0.742 0.742 0.686 0.619 $0.706(\pm 0.052)$
QUSL 0.890 0.792 0.769 0.766 0.731 0.790 (±($\pm$( ±0.059)
The experimental results show that the performance of the three networks is roughly on par, which is already an encouraging outcome. Notably, SliQ achieved comparable performance to the classical network using only about $200$ parameters, demonstrating the potential of quantum methods in parameter efficiency. QUSL outperforms SliQ in effectiveness, and also shows better performance in quantum circuit efficiency and quantum resource utilization, making it more adaptable to current quantum hardware conditions.